OpenStudy (anonymous):

If G(F(x))=x^2-3, and G(x)= x-3, find F(3)

OpenStudy (anonymous):

somebody help me please!!

OpenStudy (jdoe0001):

well is just a function composition $$g(f(x)) = x^2-3\\ g(x)= x-3 $$ so, what's f(x)? f(x) is whatever was INPUT to make "x" in x-3, to \(x^2-3\)

OpenStudy (jdoe0001):

so $$ g(x) = x-3\\ f(x) = \color{red}{x^2}\\ g(f(x)) = g(\color{red}{x^2}) = (\color{red}{x^2})-3 $$

OpenStudy (jdoe0001):

so, since your $$ f(x) = \color{red}{x^2}\\ f(-3) = \color{red}{\square?}^2 $$

OpenStudy (jdoe0001):

ohh is just 3 ahem

OpenStudy (jdoe0001):

so $$ f(x) = \color{red}{x^2}\\ f(3) = \color{red}{\square?}^2 $$

OpenStudy (anonymous):

i'm just not seeing where the x^2 comes from,

OpenStudy (jdoe0001):

g(x) = x-3 g( SOMETHING ) = ( SOMETHING ) -3 g ( f(x) ) = \(x^2 -3\)

OpenStudy (anonymous):

oh i see it now

OpenStudy (anonymous):

thank you for your help

OpenStudy (jdoe0001):

yw